Circular No. 6900

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bv IN HS 1035+4758 A. Yu. Kniazev and N. Arkhyz, Special Astrophysical Observatory

(SAO), Russian Academy of Sciences; D. Engels, Hamburger Sternwarte; S. A. Pustilnik and A. V. Ugryumov, SAO; and J. Masegosa, Institute de Astrofizica de Andalucia, Granada, report their discovery of a supernova located 3′′.1 west and 3′′.7 south of the emission-line galaxy HS 1035+4758 (α = 10^{h}38^{m}25^{s}.71, δ = +47^{o}42′36′′.5, equinox 2000.0; redshift z = 0.0050), which is a typical blue compact dwarf galaxy (size ∼ 20′′) with strong H II emission in its central part. Spectra of the new star (which is certainly absent on the Digital Sky Survey), obtained on Apr. 6.914 (range 370– 610 nm) and 19.774 UT (range 370–740 nm) with resolution ∼ 0.4 nm, are characteristic to the spectra of type-IIP supernovae in their transition phase, ∼ 1 to 2 months after peak brightness. They show broad emission lines of Hα and Hβ with redshift close to that of host halaxy, and P-Cyg absorption in Balmer series up to H_{7}. Two broad absorptions of multiplets Fe II (42) 454.9–455.5-nm and Fe II (37) 516.9-nm are well seen, blueshifted in relative systemic velocity by ∼ 4000 km/s. The FWZI of Hα is 13 500 km/s, and its centroid is shifted blueward ∼ 120 km/s (Apr. 19) relative systemic velocity of the galaxy. Broad emission Hβ is blueshifted by ∼ 200 km/s on Apr. 6, but redshifted by 300 km/s on Apr. 19. Broad and strong Balmer absorption lines have central velocities blueshifted relative to those of broad emission lines by ∼ 5500 km/s; they show a systematic trend of FWHM from 4500 km/s in Hα to 1900 km/s in Hδ. The magnitude of SN 1998bv, estimated from the calibrated spectrum of Apr. 6, is B ∼ 17.1; on Apr. 24, a CCD image yielded B = 17.4, B − V = +0.85.

XTE J1806–246 R. M. Hjellming, National Radio Astronomy Observatory (NRAO);

A. J. Mioduszewski, Joint Institute for VLBI in Europe and NRAO; and M. P. Rupen, NRAO, report: “On May 10.49 UT, a radio source was observed with the Very Large Array in the field of the x-ray transient XTE J1806–246 at α = 18^{h}06^{m}50^{s}.72 ± 0^{s}.04, δ = −24^{o}35′28′′.6 ± 0′′.8 (equinox 2000.0). This radio source is well within the error region for the RXTE PCA position reported on IAUC 6891. On May 10.49 the radio source had flux densities of 0.8 ± 0.3 and 0.8 ± 0.2 mJy at 1.43 and 4.86 GHz, respectively. It was not present on May 3 and 6, with upper limits of 0.3 mJy at the same frequencies. This variability makes the source a candidate for the radio counterpart of XTE J1806–246.”

1998 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green